# Group-in-Group Policy Optimization for LLM Agent Training

### v2

@@ 1 Introduction
Large Language Models (LLMs) [1, 2, 3, 4] have leapt from static question-answer systems to versatile agents that perceive, reason, and act in open-ended environments. For instance, they now power embodied assistants that navigate simulated homes [5, 6], mobile and web navigators that plan multi-step browsing sessions [7, 8, 9, 10], and autonomous explorers in interactive games [11, 12]. In these settings, LLM agents need to perceive, reason, and act in multi-turn loops, which requires not only language understanding but also long-horizon planning and decision-making.
Reinforcement learning (RL) [13] has become a crucial recipe for post-training LLMs, leading to frontier models like OpenAI o1 [14] and DeepSeek R1 [15]. In particular, group-based RL algorithms such as RLOO [16, 17] and GRPO [18] have proven especially effective in large-scale training. These methods replace value-function estimation with simple yet powerful relative advantage estimation within groups of rollouts. This group-based advantage computation enjoys favorable properties such as low memory overhead, critic-free optimization, and scalability to large models. However, their successes have so far been largely limited to single-turn tasks such as math problem solving [19, 20] and code generation [21], where reward arrives immediately and credit assignment is straightforward.
In contrast, LLM agents operating in external environments face fundamentally different learning landscapes. Their behavior unfolds over long episodes with tens of decision steps and tens of thousands of tokens (e.g., an ALFWorld [5] episode may include up to 50 steps and over 20k tokens). Rewards are typically sparse (sometimes arriving only at the end of an episode), and the impact of any individual action may only manifest much later in the trajectory. This substantially complicates the credit assignment for individual steps and increases the challenge of policy optimization. Naively applying existing group-based RL algorithms in such settings collapses step-level distinctions, undermining their effectiveness. Hence, these limitations raise a core question:
@@
To address this, we introduce Group-in-Group Policy Optimization (GiGPO), a new group-based RL algorithm that nests two-dimensional notions of credit assignment, better-suited for the multi-turn optimization of LLM agents. (i) At the episode level, GiGPO samples a group of complete trajectories under identical task and initial-state conditions, and computes macro relative advantages based on total returns like vanilla GRPO [18]. This captures the overall effectiveness of each trajectory and reflects the completeness of task execution. (ii) At the step level, GiGPO introduces a novel anchor state grouping mechanism for fine-grained relative advantages estimation. Specifically, it retroactively identifies the repeated environment states, termed as anchor states, across trajectories and uses them as anchors to construct step-level groups that allow for localized credit assignment.
@@
We first evaluate GiGPO on two long-horizon agent benchmarks: ALFWorld [5], which tests embodied task planning in simulated household environments, and WebShop [22], which simulates complex, goal-driven web interactions. In addition, we study multi-turn tool-integrated reasoning on search-augmented QA tasks. Our experiments with Qwen2.5-1.5B/3B/7B-Instruct [3] show that GiGPO consistently outperforms prompt-based agents, actor-critic baselines, and prior group-based RL methods. In particular, GiGPO injects fine-grained, step-level credit signals that sharpen policy learning of agents over horizons and achieves performance improvement of > 12% on ALFWorld and > 9% on WebShop over GRPO, along with remarkable performance on search-based QA tasks (42.1% on 3B and 47.2% on 7B). These gains come without compromising the core strengths of group-based RL (only < 0.002% time cost), making GiGPO a versatile and high-utility algorithm for the optimization of LLM agents.
@@ 2 Related Work
LLMs as decision-making agents.   The use of large language models (LLMs) as autonomous agents has expanded rapidly across domains such as program generation [23], smart device operation [24, 25, 26, 27], interactive gameplay [11], and robot behavior control [28]. Early works typically relied on leveraging pre-trained, frozen models through carefully designed prompting methods (like ReAct [29] and Reflexion [30]), enhanced memory and retrieval systems [12, 31], and integration with external tools [32, 33, 34]. More recent research has shifted toward adapting model parameters with supervised fine-tuning (SFT) [24] or RL [13], enabling agents to learn directly from environment interaction rather than static prompts or handcrafted workflows, which we introduce below.
@@
Reinforcement learning for LLM agents. RL has played a pivotal role in enabling LLM agents to operate in dynamic, open-ended environments. Early work applied classical RL algorithms such as DQN [35] to train LLM agents in text-based games [36] and later research [37, 38, 39, 40, 41] started to employ value-based methods, such PPO [42] and AWR [43], in more diverse and interactive agent scenarios including Android device control [44], embodied ALFWorld [5], and card games [45]. More recent approaches have extended RL training to complex web-based and application-centered tasks. For instance, ArCHer [46] and AgentQ [47] target the WebShop benchmark [22], but require intricate designs and computation overhead such as additional value networks or Monte Carlo Tree Search (MCTS) [48]. CoSo [10] introduces an entropy-based RL method that enhances the performance of agents. Going further, LOOP [49] introduces a hybrid method combining REINFORCE leave-one-out (RLOO) [16, 17] with PPO-style updates, achieving state-of-the-art results in AppWorld [50]. RAGEN [51] introduces a trajectory-level GRPO that concatenates all states, intermediate reasoning, and actions into a unified episode-level response. However, it faces scalability challenges in long-horizon tasks (e.g., in ALFWorld, which involves up to 50 steps).
@@
Reinforcement learning for large language models. An early and influential application of RL in LLMs is the Reinforcement Learning from Human Feedback (RLHF) [52, 53, 54, 55], which focuses on aligning LLMs to human preferences. Most recent works have explored using RL to enhance the capabilities of reasoning and logical deduction in LLMs [56, 15]. In particular, group-based RL algorithms have emerged as promising alternatives to traditional methods like PPO [42]. These methods, such as RLOO [16, 17], GRPO [18], Dr. GRPO [19], DAPO [20], and CPPO [57], avoid introducing extra value functions by leveraging a group of samples from the same query and estimate the advantages accordingly. This enables the large-scale RL training and has shown strong results in tasks such as mathematical reasoning [15], search [58, 59], and tool use [60, 61]. Our work is closely related to this line of research, with a focus on training LLM agents. We aim to retain the benefits of group-based RL, such as critic-free learning and efficiency, while introducing finer-grained credit assignment. Moreover, the hierarchical core of GiGPO is orthogonal to existing group-based RL approaches, making it fully compatible and capable of incorporating them to enhance performance.
@@ 3 Preliminaries
Group-based RL.   Recent RL works converge on a simple recipe for training LLMs: for a given task description $x$, the LLM samples a group of $N$ candidate trajectories $\{\bm{\tau}_{1},\bm{\tau}_{2},\dots,\bm{\tau}_{N}\}$, each corresponding to one full episode rollout under $\pi_{\theta_{\text{old}}}$. Each trajectory $\bm{\tau}_{i}$ receives a scalar reward $R(\bm{\tau}_{i})$ reflecting the overall quality or success of the generated outcome. Instead of estimating advantages using separate value functions like PPO [42], group-based RL methods compute advantages purely based on the statistics internal to the sampled group:
@@
For example, in GRPO [18], the advantage of each trajectory is estimated by normalizing its reward with respect to the group’s mean and standard deviation. This design is highly memory-efficient and can scale effectively to the large batch sizes and model sizes typical in modern LLM training, making it a practical and scalable choice for large-scale RL training.
@@ 4 Training LLM Agents with GiGPO
While group-based RL algorithms [18, 15] have proven highly effective for training LLMs in single-turn tasks, their extension to multi-step agent settings faces critical challenges in credit assignment. Figure 1 illustrates this gap. Vanilla GRPO (left) treats each trajectory as a whole and computes a single relative advantage for the entire episode, which fails to provide actionable feedback for individual steps. A natural remedy is to roll out multiple single-step actions for each state $\bm{s}_{t}$ via $\pi_{\theta_{\text{old}}}$ as shown in Figure 1 (middle). However, this approach quickly becomes impractical due to the substantial overhead of extra LLM forward passes and the difficulty of evaluating rewards for hypothetical actions never actually executed.
To overcome these challenges, we propose our Group-in-Group Policy Optimization (GiGPO) in this section. Similar to prior works [49, 51], GiGPO begins by sampling groups of trajectories under identical tasks and initial environment states. It then introduces a two-level grouping structure: preserving episode-level grouping for holistic performance comparison, while dynamically constructing an additional set of step-level groups by retroactively aggregating actions encountering the same environment states. This “group-in-group” construction yields two complementary advantages: (1) Episode relative advantages capture the holistic effectiveness of each trajectory, providing a stable, global training signal. (2) Step relative advantages zoom in on which actions outperform their peers within the same state, endowing the gradient with fine-grained credit.
@@ 4.1 Episode Relative Advantages
In GRPO [18], the default normalization factor is defined as the standard deviation, i.e., $F_{\text{norm}}=\text{std}$. However, this may introduce a difficulty bias [19], where trajectories from low-variance groups (e.g., very easy or hard tasks) receive disproportionately large gradients. In the context of the LLM agent, where tasks often involve very long horizons, this effect tends to emerge frequently, potentially affecting the stability of updates. As an alternative, we also consider a fixed normalization factor $F_{\text{norm}}=1$, which yields an unbiased Leave-One-Out estimator [16] (see Appendix B for details). This simple adjustment helps stabilize training in some challenging agent scenarios.
@@ 4.2 Step Relative Advantages
How does step-level group work?   We present an intuitive illustration in Figure 3 to show the utility of the step relative advantages. We consider two example trajectories from the set $\{\bm{\tau}_{i}\}_{i=1}^{N}$. In $\tau_{1}$, the agent first selects the 2nd Item (incorrect), then returns to the previous page and selects the 1st Item (correct), successfully completing the task. Due to temporal discounting (Equation (5)), the earlier action (2nd Item) receives a lower discounted return than the later correct one (1st Item). In $\tau_{2}$, the agent clicks the Next Page, ultimately failing to find the target and receiving no reward. By aggregating these actions into the same step-level group based on their shared anchor state, GiGPO computes their relative advantages and yields a clear preference ordering: $A^{S}(\textit{1st Item})>A^{S}(\textit{2nd Item})>A^{S}(\textit{Next Page})$. This ranking successfully captures fine-grained distinctions in long-term utility that are missed by prior group-based RL methods [17, 18, 20]. While this example illustrates GiGPO’s effectiveness in sparse-reward environments, its advantages extend naturally to dense-reward scenarios, where per-step rewards can be fully leveraged to assess the relative quality of individual actions within shared states.
@@ 4.3 Group-in-Group Policy Optimization
where $\rho_{\theta}(\bm{a}_{t}^{(i)})=\frac{\pi_{\theta}(\bm{a}_{t}^{(i)}\mid\bm{s}_{t}^{(i)},x)}{\pi_{\theta_{\text{old}}}(\bm{a}_{t}^{(i)}\mid\bm{s}_{t}^{(i)},x)}$ is the importance sampling ratio, $\beta$ controls the strength of the KL penalty encouraging proximity to a reference policy $\pi_{\text{ref}}$. We present the pseudo code in Appendix C.
@@ 5.1 Experiment Setup
Benchmarks.   We first train the LLM agents on two challenging benchmarks: ALFWorld [5] and WebShop [22]. ALFWorld is an embodied environment designed to assess the ability of LLM agents to perform multi-step decision-making. In each episode, the agent receives a text goal and must accomplish it through multi-turn interaction with the environment. It includes 3,827 task instances across six categories of common household activities: Pick & Place (Pick), Examine in Light (Look), Clean & Place (Clean), Heat & Place (Heat), Cool & Place (Cool), and Pick Two & Place (Pick2). WebShop is a complex, web-based interactive environment designed to test the LLM agents in realistic online shopping scenarios. To complete the task, the agent must interact with a simulated HTML-based shopping website to search for, navigate to, and ultimately purchase a suitable item. It contains over 1.1 million products and 12k user instructions, providing a rich and diverse action space. In addition, we also evaluate the multi-turn tool calling performance of GiGPO on search-augmented QA tasks, including single-hop QA datasets (NQ [62], TriviaQA [63], and PopQA [64]) and multi-hop QA datasets (HotpotQA [65], 2Wiki [66], MuSiQue [67], and Bamboogle [68]).
Baselines.   For ALFWorld and WebShop, we compare our approach with a range of competitive baselines: (1) Closed-source LLMs: GPT-4o [1] and Gemini-2.5-Pro [2], which represent state-of-the-art capabilities in general-purpose reasoning and language understanding. (2) Prompting agents: ReAct [29] and Reflexion [30], which rely on in-context prompting to guide multi-step behavior without parameter updates. (3) RL training methods: PPO [42], a widely used actor-critic algorithm that requires an additional value model, and group-based critic-free methods RLOO [16, 17] and GRPO [18], which perform advantage estimation over trajectory groups. For search-augmented QA, we compare GiGPO with R1-Instruct, Search-R1 [58], ZeroSearch [59], and StepSearch [69].
Training details.   We use Qwen2.5-1.5B/3B/7B-Instruct [3] as our base models. The weighting coefficient $\omega$ is set to $1$ with no further tuning. For ALFWorld and WebShop, all RL training methods (including ours and the baselines) use exactly the same hyperparameter configurations. The rollout group size $N$ for group-based RL methods is set to 8. For search-augmented QA, we follow the same settings in Search-R1 [58]. We use E5 [70] as the retriever. The rollout group size $N$ is set to 5 and the max turn is set to 4. Moreover, we incorporate similarity-based GiGPO, where anchor state grouping is performed by grouping two states if their similarity (longest matching subsequence) exceeds the threshold of 0.9. Full training settings and hyperparameter details are provided in Appendix D.1.

### v3

@@ 1 Introduction
Large Language Models (LLMs) achiam2023gpt ; team2023gemini ; yang2024qwen2 ; liu2024deepseek have leapt from static question-answer systems to versatile agents that perceive, reason, and act in open-ended environments. For instance, they now power embodied assistants that navigate simulated homes shridhar2020alfworld ; li2024embodied , mobile and web navigators that plan multi-step browsing sessions furuta2024multimodal ; zheng2024gpt ; gou2025navigating ; feng2025towards , and autonomous explorers in interactive games wang2023voyager ; wang2024mobile . In these settings, LLM agents need to perceive, reason, and act in multi-turn loops, which requires not only language understanding but also long-horizon planning and decision-making.
Reinforcement learning (RL) sutton2018reinforcement has become a crucial recipe for post-training LLMs, leading to frontier models like OpenAI o1 openai2024 and DeepSeek R1 guo2025deepseek . In particular, group-based RL algorithms such as RLOO kool2019buy ; ahmadian2024back and GRPO shao2024deepseekmath have proven especially effective in large-scale training. These methods replace value-function estimation with simple yet powerful relative advantage estimation within groups of rollouts. This group-based advantage computation enjoys favorable properties such as low memory overhead, critic-free optimization, and scalability to large models. However, their successes have so far been largely limited to single-turn tasks such as math problem solving liu2025understanding ; yu2025dapo and code generation wei2025swe , where reward arrives immediately and credit assignment is straightforward.
In contrast, LLM agents operating in external environments face fundamentally different learning landscapes. Their behavior unfolds over long episodes with tens of decision steps and tens of thousands of tokens (e.g., an ALFWorld shridhar2020alfworld episode may include up to 50 steps and over 20k tokens). Rewards are typically sparse (sometimes arriving only at the end of an episode), and the impact of any individual action may only manifest much later in the trajectory. This substantially complicates the credit assignment for individual steps and increases the challenge of policy optimization. Naively applying existing group-based RL algorithms in such settings collapses step-level distinctions, undermining their effectiveness. Hence, these limitations raise a core question:
@@
To address this, we introduce Group-in-Group Policy Optimization (GiGPO), a new group-based RL algorithm that nests two-dimensional notions of credit assignment, better-suited for the multi-turn optimization of LLM agents. (i) At the episode level, GiGPO samples a group of complete trajectories under identical task and initial-state conditions, and computes macro relative advantages based on total returns like vanilla GRPO shao2024deepseekmath . This captures the overall effectiveness of each trajectory and reflects the completeness of task execution. (ii) At the step level, GiGPO introduces a novel anchor state grouping mechanism for fine-grained relative advantages estimation. Specifically, it retroactively identifies the repeated environment states, termed as anchor states, across trajectories and uses them as anchors to construct step-level groups that allow for localized credit assignment.
@@
We first evaluate GiGPO on long-horizon agent benchmarks: ALFWorld shridhar2020alfworld , which tests embodied task planning in simulated household environments, and WebShop yao2022webshop , which simulates complex, goal-driven web interactions. In addition, we study multi-turn tool-integrated reasoning on search-augmented QA tasks. Our experiments with Qwen2.5-1.5B/3B/7B-Instruct yang2024qwen2 show that GiGPO consistently outperforms prompt-based agents, actor-critic baselines, and prior group-based RL methods. In particular, GiGPO injects fine-grained, step-level credit signals that sharpen policy learning of agents over horizons and achieves performance gains of > 12% on ALFWorld and > 9% on WebShop over GRPO, along with remarkable performance on search-based QA tasks (42.1% on 3B and 47.2% on 7B). These gains come without compromising the core strengths of group-based RL (only < 0.002% time cost), making GiGPO a versatile and high-utility algorithm for LLM agents.
@@ 2 Related Work
LLMs as decision-making agents.   The use of large language models (LLMs) as autonomous agents has expanded rapidly across domains such as program generation zhang2024codeagent , smart device operation zhang2023you ; hong2024cogagent ; gur2023real ; hu2024dawn , interactive gameplay wang2023voyager , and robot behavior control brohan2023rt . Early works typically relied on leveraging pre-trained, frozen models through carefully designed prompting methods (like ReAct yao2023react and Reflexion shinn2024reflexion ), enhanced memory and retrieval systems wang2024mobile ; tan2024cradle , and integration with external tools schick2023toolformer ; xie2024osworld ; zhang2024ufo . More recent research has shifted toward adapting model parameters with supervised fine-tuning (SFT) zhang2023you or RL sutton2018reinforcement , enabling agents to learn directly from environment interaction rather than static prompts or handcrafted workflows, which we introduce below.
@@
Reinforcement learning for LLM agents. RL has played a pivotal role in enabling LLM agents to operate in dynamic, open-ended environments. Early work applied classical RL algorithms such as DQN mnih2015human to train LLM agents in text-based games narasimhan2015language and later research tan2024true ; wen2024reinforcing ; zhai2024fine ; bai2024digirl ; wang2024distrl started to employ value-based methods, such PPO schulman2017proximal and AWR peng2019advantage , in more diverse and interactive agent scenarios including Android device control rawles2024androidinthewild , embodied ALFWorld shridhar2020alfworld , and card games brockman2016openai . More recent approaches have extended RL training to complex web-based and application-centered tasks. For instance, ArCHer zhou2024archer and AgentQ putta2024agent target the WebShop benchmark yao2022webshop , but require intricate designs and computation overhead such as additional value networks or Monte Carlo Tree Search (MCTS) silver2017mastering . CoSo feng2025towards introduces an entropy-based RL method that enhances the performance of agents. Going further, LOOP chen2025reinforcement introduces a hybrid method combining REINFORCE leave-one-out (RLOO) kool2019buy ; ahmadian2024back with PPO-style updates, achieving state-of-the-art results in AppWorld trivedi2024appworld . RAGEN wang2025ragen introduces a trajectory-level GRPO that concatenates all states, intermediate reasoning, and actions into a unified episode-level response. However, it faces scalability challenges in long-horizon tasks (e.g., in ALFWorld, which involves up to 50 steps).
@@
Reinforcement learning for large language models. An early and influential application of RL in LLMs is the Reinforcement Learning from Human Feedback (RLHF) ziegler2019fine ; stiennon2020learning ; ouyang2022training ; rafailov2024direct , which focuses on aligning LLMs to human preferences. Most recent works have explored using RL to enhance the capabilities of reasoning and logical deduction in LLMs team2025kimi ; guo2025deepseek . In particular, group-based RL algorithms have emerged as promising alternatives to traditional methods like PPO schulman2017proximal . These methods, such as RLOO kool2019buy ; ahmadian2024back , GRPO shao2024deepseekmath , Dr. GRPO liu2025understanding , DAPO yu2025dapo , and CPPO lin2025cppo , avoid introducing extra value functions by leveraging a group of samples from the same query and estimate the advantages accordingly. This enables the large-scale RL training and has shown strong results in tasks such as mathematical reasoning guo2025deepseek , search jin2025search ; sun2025zerosearch , and tool use qian2025toolrl ; wang2025otc . Our work is closely related to this line of research, with a focus on training LLM agents. We aim to retain the benefits of group-based RL, such as critic-free learning and efficiency, while introducing finer-grained credit assignment. Moreover, the hierarchical core of GiGPO is orthogonal to existing group-based RL approaches, making it fully compatible and capable of incorporating them to enhance performance.
@@ 3 Preliminaries
Group-based RL.   Recent RL works converge on a simple recipe for training LLMs: for a given task description $x$, the LLM samples a group of $N$ candidate trajectories $\{\bm{\tau}_{1},\bm{\tau}_{2},\dots,\bm{\tau}_{N}\}$, each corresponding to one full episode rollout under $\pi_{\theta_{\text{old}}}$. Each trajectory $\bm{\tau}_{i}$ receives a scalar reward $R(\bm{\tau}_{i})$ reflecting the overall quality or success of the generated outcome. Instead of estimating advantages using separate value functions like PPO schulman2017proximal , group-based RL methods compute advantages purely based on the statistics internal to the sampled group:
@@
For example, in GRPO shao2024deepseekmath , the advantage of each trajectory is estimated by normalizing its reward with respect to the group’s mean and standard deviation. This design is highly memory-efficient and can scale effectively to the large batch sizes and model sizes typical in modern LLM training, making it a practical and scalable choice for large-scale RL training.
@@ 4 Training LLM Agents with GiGPO
While group-based RL algorithms shao2024deepseekmath ; guo2025deepseek have proven highly effective for training LLMs in single-turn tasks, their extension to multi-step agent settings faces critical challenges in credit assignment. Figure 1 illustrates this gap. Vanilla GRPO (left) treats each trajectory as a whole and computes a single relative advantage for the entire episode, which fails to provide actionable feedback for individual steps. A natural remedy is to roll out multiple single-step actions for each state $\bm{s}_{t}$ via $\pi_{\theta_{\text{old}}}$ as shown in Figure 1 (middle). However, this approach quickly becomes impractical due to the substantial overhead of extra LLM forward passes and the difficulty of evaluating rewards for hypothetical actions never actually executed.
To overcome these challenges, we propose our Group-in-Group Policy Optimization (GiGPO) in this section. Similar to prior works chen2025reinforcement ; wang2025ragen , GiGPO begins by sampling groups of trajectories under identical tasks and initial environment states. It then introduces a two-level grouping structure: preserving episode-level grouping for holistic performance comparison, while dynamically constructing an additional set of step-level groups by retroactively aggregating actions encountering the same environment states. This “group-in-group” construction yields two complementary advantages: (1) Episode relative advantages capture the holistic effectiveness of each trajectory, providing a stable, global training signal. (2) Step relative advantages zoom in on which actions outperform their peers within the same state, endowing the gradient with fine-grained credit.
@@ 4.1 Episode Relative Advantages
In GRPO shao2024deepseekmath , the default normalization factor is defined as the standard deviation, i.e., $F_{\text{norm}}=\text{std}$. However, this may introduce a difficulty bias liu2025understanding , where trajectories from low-variance groups (e.g., very easy or hard tasks) receive disproportionately large gradients. In the context of the LLM agent, where tasks often involve very long horizons, this effect tends to emerge frequently, potentially affecting the stability of updates. As an alternative, we also consider a fixed normalization factor $F_{\text{norm}}=1$, which yields an unbiased Leave-One-Out estimator kool2019buy (see Appendix C for details). This simple adjustment helps stabilize training in some challenging agent scenarios.
@@ 4.2 Step Relative Advantages
How does step-level group work?   We present an intuitive illustration in Figure 3 to show the utility of the step relative advantages. We consider two example trajectories from the set $\{\bm{\tau}_{i}\}_{i=1}^{N}$. In $\tau_{1}$, the agent first selects the 2nd Item (incorrect), then returns to the previous page and selects the 1st Item (correct), successfully completing the task. Due to temporal discounting (Equation (5)), the earlier action (2nd Item) receives a lower discounted return than the later correct one (1st Item). In $\tau_{2}$, the agent clicks the Next Page, ultimately failing to find the target and receiving no reward. By aggregating these actions into the same step-level group based on their shared anchor state, GiGPO computes their relative advantages and yields a clear preference ordering: $A^{S}(\textit{1st Item})>A^{S}(\textit{2nd Item})>A^{S}(\textit{Next Page})$. This ranking successfully captures fine-grained distinctions in long-term utility that are missed by prior group-based RL methods ahmadian2024back ; shao2024deepseekmath ; yu2025dapo . While this example illustrates GiGPO’s effectiveness in sparse-reward environments, its advantages extend naturally to dense-reward scenarios, where per-step rewards can be fully leveraged to assess the relative quality of individual actions within shared states.
@@ 4.3 Group-in-Group Policy Optimization
where $\rho_{\theta}(\bm{a}_{t}^{(i)})=\frac{\pi_{\theta}(\bm{a}_{t}^{(i)}\mid\bm{s}_{t}^{(i)},x)}{\pi_{\theta_{\text{old}}}(\bm{a}_{t}^{(i)}\mid\bm{s}_{t}^{(i)},x)}$ is the importance sampling ratio, $\beta$ controls the strength of the KL penalty encouraging proximity to a reference policy $\pi_{\text{ref}}$. We present the pseudo code in Appendix D.
@@ 5.1 Experiment Setup
Benchmarks.   We first train the LLM agents on two challenging benchmarks: ALFWorld shridhar2020alfworld and WebShop yao2022webshop . ALFWorld is an embodied environment designed to assess the ability of LLM agents to perform multi-step decision-making. In each episode, the agent receives a text goal and must accomplish it through multi-turn interaction with the environment. It includes 3,827 task instances across six categories of common household activities: Pick & Place (Pick), Examine in Light (Look), Clean & Place (Clean), Heat & Place (Heat), Cool & Place (Cool), and Pick Two & Place (Pick2). WebShop is a complex, web-based interactive environment designed to test the LLM agents in realistic online shopping scenarios. To complete the task, the agent must interact with a simulated HTML-based shopping website to search for, navigate to, and ultimately purchase a suitable item. It contains over 1.1 million products and 12k user instructions, providing a rich and diverse action space. In addition, we also evaluate the multi-turn tool calling performance of GiGPO on search-augmented QA tasks, including single-hop QA datasets (NQ kwiatkowski2019natural , TriviaQA joshi2017triviaqa , and PopQA mallen2022not ) and multi-hop QA datasets (HotpotQA yang2018hotpotqa , 2Wiki ho2020constructing , MuSiQue trivedi2022musique , and Bamboogle press2022measuring ).
Baselines.   For ALFWorld and WebShop, we compare our approach with a range of competitive baselines: (1) Closed-source LLMs: GPT-4o achiam2023gpt and Gemini-2.5-Pro team2023gemini , which represent state-of-the-art capabilities in general-purpose reasoning and language understanding. (2) Prompting agents: ReAct yao2023react and Reflexion shinn2024reflexion , which rely on in-context prompting to guide multi-step behavior without parameter updates. (3) RL training methods: PPO schulman2017proximal , a widely used actor-critic algorithm that requires an additional value model, and group-based critic-free methods RLOO kool2019buy ; ahmadian2024back and GRPO shao2024deepseekmath , which perform advantage estimation over trajectory groups. For search-augmented QA, we compare GiGPO with R1-Instruct, Search-R1 jin2025search , ZeroSearch sun2025zerosearch , and StepSearch wang2025stepsearch .
Training details.   We use Qwen2.5-1.5B/3B/7B-Instruct yang2024qwen2 as our base models. The weighting coefficient $\omega$ is set to $1$ with no further tuning. For ALFWorld and WebShop, all RL training methods (including ours and the baselines) use exactly the same hyperparameter configurations. The rollout group size $N$ for group-based RL methods is set to 8. For search-augmented QA, we follow the same settings in Search-R1 jin2025search . We use E5 wang2022text as the retriever. The rollout group size $N$ is set to 5 and the max turn is set to 4. Moreover, we incorporate similarity-based GiGPO, where anchor state grouping is performed by grouping two states if their similarity (longest matching subsequence) exceeds the threshold of 0.9. Full training settings and hyperparameter details are provided in Appendix E.1.